OpenStudy (anonymous):

ok i have -1/2 -(-3) ---------- -3 - 5 then i get lost because of the fraction

OpenStudy (anonymous):

solve the top, what is -.5 - (-3)

OpenStudy (anonymous):

-0.5+3=2.5 -3-5=-8 2.5/-8 -5/16

OpenStudy (anonymous):

also you can ask your questions here anna instead of making another question place if its a follow up on your original question you asked